OpenStudy (anonymous):

5/8q^6 multiplied by 4q^5/3 is what?

OpenStudy (anonymous):

=(5/3)(4q^5/8q^6) =(5/3)(q^5/2q^6) =(5/3)(q^-1/2) =(5/3)(1/2q)(Since negative exponents go to denominator) =5/6q

OpenStudy (anonymous):

Thank you!